# DeepASA: An Object-Oriented Multi-Purpose Network for Auditory Scene Analysis

### v2

@@ 1 Introduction
However, unlike the human auditory system, these task-specific models typically lack the capacity for relational reasoning across tasks and cues. As a result, they often fail when critical auditory cues are missing or degraded he2021deep . Recent work highlights the benefit of combining multiple auditory cues. For instance, target sound extraction (TSE), which uses class veluri2023real ; delcroix2022soundbeam , timestamp wang2022improving ; kim2024improving , and spatial cues jenrungrot2020cone ; gu2024rezero , outperforms universal sound separation (USS) in challenging scenarios kong2023universal ; zmolikova2023neural . Similarly, joint SED and DoAE in sound event localization and detection (SELD) has been shown to enhance performance by leveraging interdependent auditory information adavanne2018sound ; he2021sounddet .
@@
1. Object-Oriented Processing (OOP): We design an encoder-decoder framework that isolates features of individual sound objects. The encoder incorporates a dynamic short-time Fourier transform (STFT) module and a transformer-based feature aggregator. These are followed by an object feature separator that disentangles per-object auditory representations. These object-centric features are then passed to multiple sub-decoders, each responsible for estimating a specific auditory parameter (e.g., class, DoA, timestamp) without requiring manual association or permutation invariant training between estimated auditory parameters.
@@ 2.1 Overall framework: DeepASA
We aim to estimate the auditory information of up to $J$ foreground sources, including classes, timestamps of onsets and offsets, DoA trajectories, and multichannel waveforms of the direct and reverb audio signals, as well as one multichannel noise signal, from the multichannel audio mixture.
@@
As illustrated in Figure 2, the overall architecture of DeepASA comprises three components: audio encoder, object separator, and sub-decoders. The audio encoder extracts basic features from the $M$-channel input mixture. The object separator separates audio features extracted from the audio encoder into $J+1$ object features, each of which encapsulates the auditory information of a single sound object. The sub-decoders then estimate the auditory information from each object feature.
@@ 2.2 Audio encoder
Dynamic STFT with time-varying learnable window The audio encoder of DeepASA incorporates our proposed architecture, Dynamic STFT, for adaptive temporal focusing on the input waveform. Conventional STFT uses a fixed window at all times, yet the locations and durations of salient information within time frames can vary across the waveform. Accordingly, we employ a learnable Gaussian window whose mean $(\mu_{t})$ and standard deviation $(\sigma_{t})$ are predicted at every frame, forming $\boldsymbol{\mu}\in\mathbb{R}^{T}$ and $\boldsymbol{\sigma}\in\mathbb{R}^{T}$ vectors for the total number of time frames $T$. The frame-wise mean $(\mu_{t})$ aligns the center of the window on the most informative region, while $(\sigma_{t})$ sculpts its taper: a large $\sigma_{t}$ flattens the window toward a rectangle, sharpening spectral focus with a thin main lobe, whereas a small $\sigma_{t}$ contracts it toward an impulse, enhancing temporal focus.
@@
Processing STFT with the time-varying learnable window, the multichannel waveform is converted into a complex spectrogram $\mathbf{X}\in\mathbb{R}^{2M\times T\times F}$ where $F$ is the number of frequency bins. These complex spectrograms are then fed directly into an Up-Conv block composed of a 2D convolution layer followed by layer normalization, to increase the channel dimension from $2M$ to $C$ while preserving the temporal and spectral dimensions.
@@ 2.4 Sub-decoders
The sub-decoders decode each object feature into the direct, reverberant, and noise audio signals, class, timestamp, and DoA outputs. The detailed architectures of sub-decoders are depicted in Figure 4(b).
@@
SED decoder The SED decoder decodes the object features into three different outputs related to the SELD task: a predicted class probability, a single-channel object activation (onset and offset) curve, and a sound event map presenting activation of individual classes across time. Ideally, a single object is only mapped to a single class label and activation curve, and the sound event map would be redundant. However, the object features are not perfectly separated in practice, so training with the auxiliary sound event map was useful for improving the SELD performance. During evaluation, only the predicted class probability and object activation curve are used to calculate the SELD scores. The SED decoder combines a pre-trained audio teacher-student transformer (ATST) shao2024fine with two separate convolutional recurrent neural network (CRNN) cakir2017convolutional -based branches summarizing time and frequency information. We employ the pre-trained ATST to utilize various features learned from AudioSet, a large-scale dataset containing a wide range of audio classes gemmeke2017audio . The features from ATST are combined with auxiliary features developed from another Conv2d layer. The concatenated features are analyzed in time by Time (T)-GRU and summarized to predict the sound event map. For predicting a single-channel activation curve, adaptive statistical pooling (ASP) okabe2018attentive and one fully connected layer are applied along the class dimension. The class probability is predicted by superposing the logit derived from Frequency (F)-CRNN, which prioritizes the spectral information using Conv2d layers and F-GRU, with the logit obtained from ASP. For inactive sources, the class decoder is trained to predict an equal probability of $1/C$ across all $C$ classes, while the sound event and timestamp decoders are trained to produce zero values for all time frames. The detailed specifications of the SED decoder are provided in Appendix A.
@@ 2.5 Chain-of-inference
Temporal coherence matching In TCM, temporal coherence between SED and DoAE is evaluated through the multi-clue attention mechanism. One branch of multi-clue attention takes SED as queries and DoA as keys and values, and the attention output is added to SED to refine the SED information. Conversely, the other multi-clue attention branch uses DoA as queries and SED as keys and values. The outputs of the two multi-clue attentions are passed through two individual linear layers to produce the $D$-channel clues and then superposed for fusion. The fused clue is interpolated by a transposed convolution layer to match its time dimension $T^{\prime}$ with that of object features ($T$).
@@
For efficient training of the CoI architecture, we utilize multi-stage training. That is, Net 1 of Figure 5 is trained first, then Net 2 is trained with Net 1 frozen.
@@ 3.1 Datasets
The proposed architecture is trained and evaluated on three spatial audio datasets: ASA2, MC-FUSS, and STARSS23. Auditory Scene Analysis dataset V2 (ASA2) provides ground truth signals and labels for all acoustic parameters addressed in this study, including comprehensive auditory information for moving source separation and scene analysis in noisy, reverberant environment. Compared to a previously proposed dataset lee2025deft , ASA2 provides a direct and reverberant audio signal, and the maximum number of foreground sources has increased from four to five. The balance in the number of audio clips has been changed to be proportional to the number of sources constituting the scene. Multichannel-Free Universal Sound Separation (MC-FUSS) aizawa2023unsupervised is a benchmark dataset for multichannel USS, including mixtures of 1–4 sound sources, with one being the background source. STARSS23 shimada2023starss23 is a dataset for the SELD task, addressing various sound events with 3.8 hours of audio-video data collected from real environments. This dataset generally requires data augmentation, so additional data augmentation of 20 hours was conducted using a spatial scaper roman2024spatial . All datasets were set to a 16 kHz sampling rate and a 4-second duration.
@@ 3.2 Training and model configuration
The direct and reverb decoders in the audio decoder were trained with the SA-SDR loss von2022sa . We configured a reference channel among multichannel signals and applied different weights (1:0.1) to the losses of the reference and the other channels. The noise decoder was trained by the SI-SDR loss le2019sdr . The joint loss was calculated by summing the losses with weights 1:1:0.01 for direct, reverb, and noise decoders. The losses for the SED decoder were cross-entropy, binary cross-entropy, and mean-square-error (MSE), for the class decoder, timestamp decoder, and sound event decoder, respectively. The MSE loss was used for the DoA decoder, which was summed with all aforementioned losses with equal weights. In evaluation, SI-SDRi (dB) and SDRi (dB) were used for USS performance evaluation, while SELD metrics (error rate (ER,%), F1-score (F1,%), localization error (LE, degrees), localization recall (LR, %)) adavanne2018sound were employed for SED and DoAE. We measured the model complexity by the parameter size and computational complexity.
4 Results
4.1 Ablation study of DeepASA blocks
@@
Object
separator
(-) Unfold 10.3 11.4 39.8 60.5 22.2 66.1 0.314 7.2 M 88.0 G
@@
Audio
decoder
(+) Noise decoder 11.0 11.7 30.3 69.8 21.2 76.0 0.241 8.1 (+96.8) M 98.9 G
@@
Dynamic
STFT
(+) Time-invariant window 10.7 11.4 30.2 69.8 19.3 76.0 0.238 8.1 (+96.8) M 99.0 G
@@
Table 2: Ablation study for the chain-of-inference architecture
Model variation USS SED DoAE SELD $\downarrow$ Complexities
without
audio decoder 2
(+) SED branch 11.0 11.8 26.6 71.8 18.2 76.5 0.221 10.3 (+96.8) M 101.6 G
(+) DoA branch 11.0 11.8 28.2 70.6 17.6 76.0 0.228 10.3 (+96.8) M 101.6 G
(+) SED & DoA 11.0 11.7 26.5 73.0 17.3 77.2 0.214 12.1 (+96.8) M 104.0 G
[…]
@@ 4.3 Generalization to spatial audio benchmarks & comparison with SOTA models
Model J=2 J=3 J=4 Total Param. MAC/s
ByteDance-uss kong2023universal 14.8 14.4 12.7 14.0 28.0 (+80.7) M 40.1 G
@@
DeFT-Mamba lee2025deft 18.4 17.1 13.8 16.4 4.2 M 58.2 G
@@
SELD performance comparison on STARSS23 dataset We evaluate two versions of DeepASA (with and without CoI) against SOTA SELD models, including top-ranking submissions from the DCASE 2023 challenge. DeepASA achieves the highest overall performance on the STARSS23 dataset, even without employing the audio decoder. Both versions of DeepASA surpass the performance of the top-ranked model Du_NERCSLIP_task3_report , which relied on class-dependent separation networks and model ensembling. It is worth noting that DeepASA was pre-trained on the ASA2 dataset prior to fine-tuning. To control for the effect of pre-training, we further compared DeepASA to publicly available models that were also trained using ASA2. As detailed in Appendix F, DeepASA consistently achieves SOTA performance across all evaluated tasks under the same training conditions. These results collectively demonstrate the strength of DeepASA across diverse ASA tasks and datasets, validating the effectiveness of combining class, timestamp, and DoA trajectory information through the object-oriented processing (OOP) and chain-of-inference (CoI) architecture.
@@ 5 Related Works
Universal Sound Separation (USS) USS addresses the challenge of separating various sound sources in an auditory scene where multiple sources overlap kavalerov2019universal ; wisdom2021s . Humans perform source separation by leveraging various characteristics of individual sources, such as class, timestamp, and DoA. However, existing USS methods rely on time-frequency representations, which may limit their ability to exploit such source-specific attributes kong2023universal ; gu2024rezero . To address this, the proposed method draws inspiration from human segregation processing and integrates multiple audio cues within a unified framework, resulting in improved performance on the USS task.
Target Sound Extraction (TSE) TSE refers to the separation of specific sound sources from audio mixture, with a focus on the auditory information of the source veluri2023real ; delcroix2022soundbeam . TSE plays a significant role in the source separation area, where spatial and temporal characteristics are leveraged to extract the target sound wang2022improving ; kong2023universal . A limitation of TSE is that humans must explicitly provide clues such as class, timestamp, and DoA. However, the network itself needs to iteratively infer these clues and use them to refine the separation process. To overcome this limitation, the proposed method introduces a strategy in which the clues inferred by the network are integrated to improve separation performance.

### v3

@@ 1 Introduction
However, unlike the human auditory system, these task-specific models typically lack the capacity for relational reasoning across tasks and cues. As a result, they often fail when critical auditory cues are missing or degraded he2021deep. Recent work highlights the benefit of combining multiple auditory cues. For instance, target sound extraction (TSE), which uses class veluri2023real; delcroix2022soundbeam, activation wang2022improving; kim2024improving, and spatial cues jenrungrot2020cone; gu2024rezero, outperforms universal sound separation (USS) in challenging scenarios kong2023universal; zmolikova2023neural. Similarly, joint SED and DoAE in sound event localization and detection (SELD) has been shown to enhance performance by leveraging interdependent auditory information adavanne2018sound; he2021sounddet.
@@
1. Object-Oriented Processing (OOP): We design an encoder-decoder framework that isolates features of individual sound objects. The encoder incorporates a dynamic short-time Fourier transform (STFT) module and a transformer-based feature aggregator. These are followed by an object feature separator that disentangles per-object auditory representations. These object-centric features are then passed to multiple sub-decoders, each responsible for estimating a specific auditory parameter (e.g., class, DoA, activation) without requiring manual association or permutation invariant training between estimated auditory parameters.
@@ 2.1 Overall framework: DeepASA
We aim to estimate the auditory information of up to $J$ foreground sources, including classes, activations of onsets and offsets, DoA trajectories, and multichannel waveforms of the direct and reverb audio signals, as well as one multichannel noise signal, from the multichannel audio mixture.
@@
As illustrated in Figure 2, the overall architecture of DeepASA comprises three components: audio encoder, object separator, and sub-decoders. The audio encoder extracts basic features from the $M$-channel input mixture. The object separator separates audio features extracted from the audio encoder into $J+1$ object features, each of which encapsulates the auditory information of a single sound object. The sub-decoders then estimate the auditory information from each object feature. $D,T^{\prime}$ and $C$ are denoted as channel dimension, time frames and number of classes.
@@ 2.2 Audio encoder
Dynamic STFT with time-varying learnable window The audio encoder of DeepASA incorporates our proposed architecture, Dynamic STFT, for adaptive temporal focusing on the input waveform. Conventional STFT uses a fixed analysis window at all times, yet the locations and durations of salient information within time frames can vary across the waveform. Accordingly, we employ a learnable Gaussian window whose mean $(\mu_{t})$ and standard deviation $(\sigma_{t})$ are predicted at every frame, forming $\boldsymbol{\mu}\in\mathbb{R}^{T}$ and $\boldsymbol{\sigma}\in\mathbb{R}^{T}$ vectors for the total number of time frames $T$. The frame-wise mean $(\mu_{t})$ aligns the center of the analysis window on the most informative region, while $(\sigma_{t})$ sculpts its taper: a large $\sigma_{t}$ flattens the analysis window toward a rectangle, sharpening spectral focus with a thin main lobe, whereas a small $\sigma_{t}$ contracts it toward an impulse, enhancing temporal focus.
@@
Processing STFT with the time-varying learnable window, the multichannel waveform is converted into a complex spectrogram $\mathbf{X}\in\mathbb{R}^{2M\times T\times F}$ where $F$ is the number of frequency bins. These complex spectrograms are then fed directly into an Up-Conv block composed of a 2D convolution layer followed by layer normalization, to increase the channel dimension from $2M$ to $D$ while preserving the temporal and spectral dimensions. For training a learnable window, the parameters are frozen at the beginning of the training, and when the model has converged, they are unfrozen and trained together.
@@ 2.4 Sub-decoders
The sub-decoders decode each object feature into the direct, reverberant, and noise audio signals, class, activation, and DoA outputs. The detailed architectures of sub-decoders are depicted in Figure 4(b).
@@
SED decoder The SED decoder decodes the object features of each sound source into the predicted class probability ($1\times C$), the binary object activation curve for event onset and offset ($T’\times 1$), and the predicted SED map ($T’\times C$). Here, $T$’ and $C$ denote the number of time frames and the number of classes, respectively. The predicted SED map is only used in the training stage to guide the joint detection of a sound event and class. In the inference stage, we use the SED map produced by multiplying the predicted class and activation curve, resulting in a single SED map for each sound source. This is to exclude the possibility of more than one class being mapped to the SED map for each sound object. The SED decoder combines a pre-trained audio teacher-student transformer (ATST) shao2024fine with two separate convolutional recurrent neural network (CRNN) cakir2017convolutional-based branches summarizing time and frequency information. CRNN consists of multiple repetitions of Conv2d and Maxpool2d (denoted as CNN), followed by GRU and FC layers. Specifically, among the three red branches of the SED decoder, the middle path with seven Conv+Maxpool layers outputting ($J\times T^{\prime}\times D$) is the T-CRNN path, and the bottom path with two Conv+Maxpool layers outputting ($J\times F^{\prime}\times D$) corresponds to the F-CRNN path. We employ the pre-trained ATST to utilize various features learned from AudioSet, a large-scale dataset containing a wide range of audio classes gemmeke2017audio. The features from ATST are combined with auxiliary features developed from another Conv2d layer. The concatenated features are analyzed in time by Time (T)-GRU and summarized to predict the sound event map. For predicting a single-channel activation curve, adaptive statistical pooling (ASP) okabe2018attentive and one fully connected layer are applied along the class dimension. The class probability is predicted by superposing the logit derived from Frequency (F)-CRNN, which prioritizes the spectral information using Conv2d layers and F-GRU, with the logit obtained from ASP. For inactive sources, the class decoder is trained to predict an equal probability of $1/C$ across all $C$ classes, while the sound event and activation decoders are trained to produce zero values for all time frames. The detailed specifications of the SED decoder are provided in Appendix A.
@@ 2.5 Chain-of-inference
Temporal coherence matching In TCM, temporal coherence between SED and DoAE is evaluated through the multi-clue attention mechanism. One branch of multi-clue attention takes SED as queries and DoA as keys and values, and the attention output is added to SED to refine the SED information. Conversely, the other multi-clue attention branch uses DoA as queries and SED as keys and values. The SED decoder outputs three components, but ultimately, the SED map of $T^{\prime}\times C$ is constructed for each object from the predicted class probability ($1\times C$) and object activation ($T^{\prime}\times 1$). DoA decoder also generates T’x3 output for each object. The alignment between the SEDs and DoAs is accomplished in TCM through the cross-attention by assigning the time dimension as a sequence dimension. The outputs of the two multi-clue attentions are passed through two individual linear layers to produce the $D$-channel clues and then superposed for fusion. The fused clue is interpolated by a transposed convolution layer to match its time dimension $T^{\prime}$ with that of object features ($T$).
@@
For efficient training of the CoI architecture, we utilize multi-stage training. That is, Net 1 of Figure 5 is trained first, then Net 2 is trained with up to the third DeFT-Mamba block of Net 1 frozen.
@@ 3.1 Datasets
The proposed architecture is trained and evaluated on three spatial audio datasets: ASA2, MC-FUSS, and STARSS23. Auditory Scene Analysis dataset V2 (ASA2) provides ground truth signals and labels for all acoustic parameters addressed in this study, including comprehensive auditory information for moving source separation and scene analysis in noisy, reverberant environment. ASA2 is based on the ASA dataset lee2025deft, but we modified it extensively to enable the comprehensive tasks discussed in this study. ASA2 supports (1) the separation of direct and reverberation components (for dereverberation), (2) MIMO separation, as compared to MISO separation of the ASA dataset, and (3) an increased maximum number of foreground sound sources, from four to five. The balance in the number of audio clips has been changed to be proportional to the number of sources constituting the scene. Multichannel-Free Universal Sound Separation (MC-FUSS) aizawa2023unsupervised is a benchmark dataset for multichannel USS, including mixtures of 1–4 sound sources, with one being the background source. For a fair comparison on the MC-FUSS dataset, we trained DeepASA both from scratch and with fine-tuning of the pre-trained model on the ASA 2 dataset. Also, we compared the effect of pre-trained ATST in the MC-FUSS dataset. STARSS23 shimada2023starss23 is a dataset for the SELD task, addressing various sound events with 3.8 hours of audio-video data collected from real environments. This dataset generally requires data augmentation, so additional data augmentation of 80 hours was conducted using a spatial scaper roman2024spatial. In this experiment, other SELD models were also trained using the large auxiliary datasets. The DCASE 2023 challenge allowed the use of external datasets and pre-trained models, so we applied this rule. All datasets were set to a 16 kHz sampling rate and a 4-second duration.
@@ 3.2 Training and model configuration
The direct and reverb decoders in the audio decoder were trained with the source-aggregated signal-to-distortion ratio (SA-SDR) loss von2022sa with permutation invariant training (PIT) yu2017permutation. We configured a reference channel among multichannel signals and applied different weights (1:0.1) to the losses of the reference and the other channels. The noise decoder was trained by the scale-invariant signal-to-distortion ratio (SI-SDR) loss le2019sdr. We employed one noise decoder, so the noise decoder was not trained with PIT. The joint loss was calculated by summing the losses with weights 1:1:0.01 for direct, reverb, and noise decoders. The losses for the SED decoder were cross-entropy, binary cross-entropy, and mean-square-error (MSE), for the class decoder, activation decoder, and sound event decoder, respectively. The MSE loss was used for the DoA decoder, which was summed with all aforementioned losses with equal weights. In evaluation, scale-invariant signal-to-distortion ratio improvement (SI-SDRi) (dB) and signal-to-distortion ratio improvement (SDRi) (dB) were used for USS performance evaluation, while SELD metrics (error rate (ER,%), F1-score (F1,%), localization error (LE, degrees), localization recall (LR, %)) adavanne2018sound were employed for SED and DoAE. We measured the model complexity by the parameter size and computational complexity.
@@
Object separator (-) Unfold 10.3 11.4 39.8 60.5 22.2 66.1 0.314 7.2 M 88.0 G
@@
Audio decoder (+) Noise decoder 11.0 11.7 30.3 69.8 21.2 76.0 0.241 8.1 (+96.8) M 98.9 G
@@
Dynamic STFT (+) Time-invariant window 10.7 11.4 30.2 69.8 19.3 76.0 0.238 8.1 (+96.8) M 99.0 G
@@
4 Results
4.1 Ablation study of DeepASA blocks
@@ 4.2 Ablation study of chain-of-inference (CoI)
Table 2: Ablation study for the chain-of-inference architecture
Model variation USS SED DoAE SELD $\downarrow$ Complexities
without audio decoder 2 (+) SED branch 11.0 11.8 26.6 71.8 18.2 76.5 0.221 10.3 (+96.8) M 101.6 G
(+) DoA branch 11.0 11.8 28.2 70.6 17.6 76.0 0.228 10.3 (+96.8) M 101.6 G
(+) SED & DoA 11.0 11.7 26.5 73.0 17.3 77.2 0.214 12.1 (+96.8) M 104.0 G
[…]
@@ 4.3 Generalization to spatial audio benchmarks & comparison with SOTA models
Training Model J=2 J=3 J=4 Total Param. MAC/s
From scratch ByteDance-uss kong2023universal 14.8 14.4 12.7 14.0 28.0 (+80.7) M 40.1 G
@@
DeFT-Mamba lee2025deft 18.4 17.1 13.8 16.4 3.6 M 83.8 G
@@
Fine-tuning (pre-trained on ASA2) DeFT-Mamba lee2025deft 18.4 17.1 14.1 16.6 3.6 M 83.8 G
[…]
SELD performance comparison on STARSS23 dataset We evaluate two versions of DeepASA (with and without CoI) against SOTA SELD models, including top-ranking submissions from the DCASE 2023 challenge. DeepASA achieves the highest overall performance on the STARSS23 dataset, even without employing the audio decoder. Both versions of DeepASA surpass the performance of the top-ranked model Du_NERCSLIP_task3_report, which relied on class-dependent separation networks and model ensembling. It is worth noting that DeepASA was pre-trained on the ASA2 dataset prior to fine-tuning. To control for the effect of pre-training, we further compared DeepASA to publicly available models that were also trained using ASA2. As detailed in Appendix F, DeepASA consistently achieves SOTA performance across all evaluated tasks under the same training conditions. These results collectively demonstrate the strength of DeepASA across diverse ASA tasks and datasets, validating the effectiveness of combining class, activation, and DoA trajectory information through the object-oriented processing (OOP) and chain-of-inference (CoI) architecture.
@@ 5 Related Works
Universal Sound Separation (USS) USS addresses the challenge of separating various sound sources in an auditory scene where multiple sources overlap kavalerov2019universal; wisdom2021s. Humans perform source separation by leveraging various characteristics of individual sources, such as class, activation, and DoA. However, existing USS methods rely on time-frequency representations, which may limit their ability to exploit such source-specific attributes kong2023universal; gu2024rezero. To address this, the proposed method draws inspiration from human segregation processing and integrates multiple audio cues within a unified framework, resulting in improved performance on the USS task.
Target Sound Extraction (TSE) TSE refers to the separation of specific sound sources from audio mixture, with a focus on the auditory information of the source veluri2023real; delcroix2022soundbeam. TSE plays a significant role in the source separation area, where spatial and temporal characteristics are leveraged to extract the target sound wang2022improving; kong2023universal. A limitation of TSE is that humans must explicitly provide clues such as class, activation, and DoA. However, the network itself needs to iteratively infer these clues and use them to refine the separation process. To overcome this limitation, the proposed method introduces a strategy in which the clues inferred by the network are integrated to improve separation performance.